A beaker contains 200 g of water. The water equivalent of beaker is 20 gm. The initial temperature of water in the beaker is 20°C. If 440 g of hot water at 92°C is poured in, then the final temperature, will be nearest to: (neglect heat loss)
(1) 58°C (2) 68°C (3) 73°C (4) none

Grade:10

1 Answers

Abhinav
8 years ago
knowing that no heat is lost

using Q=ms\bigtriangleup t

we can write

(200+20)\times s\times (x-20) +440\times s\times (x-92)=0

wherexis final temperature

so,x=68oC